OpenStudy (anonymous):

One force of 10 N points north while the other of 10 N points east. The total force is?

OpenStudy (anonymous):

Since North and East create a right angle you can use the pythagorean theorem. 10^2 + 10^2 = x^2 100 + 100 = x^2 200 = x^2 x = sqrt(200) which simplifies to... x = sqrt(100 * 2) x = 10*sqrt(2)

OpenStudy (anonymous):

14 N NE